# Benchmarking LLMs via Uncertainty Quantification

### v1

@@ C.1 Detailed Results of LAC and APS
Table 1 presents the average results derived from the two conformal score functions, namely, LAC and APS. In this part, we analyze the results associated with each conformal score function. The detailed results for LAC and APS are reported in Table 3 and Table 4, respectively.
It is evident that for both conformal score functions, the ranking of LLMs based on accuracy can be different from that based on uncertainty. These results reaffirm the importance of considering uncertainty in order to evaluate the performance of LLMs in a more holistic manner. Another notable observation is the difference in the uncertainty estimations produced by LAC and APS. In general, APS tends to produce larger prediction sets. More importantly, APS can lead to a significantly different ranking of LLMs based on uncertainty compared to LAC. For example, on the CI task, Qwen-14B secures the lowest uncertainty when LAC is utilized as the conformal score function. However, when APS is employed as the conformal score function, Qwen-14B is ranked fifth. This observation suggests that it is essential to average the results of the two conformal score functions to provide a more accurate quantification of uncertainty. Last but not least, both conformal score functions can achieve high coverage rates, verifying again the rationality of relying on prediction set size to estimate uncertainty.
C.2 Effects of Model Scale (Cont.)
Figures 7-10 illustrate the performance outcomes of the Llama-2 series, Yi series, DeepSeek series, and Falcon series. It is observed that while in general, increasing model size can lead to stronger performance, on some tasks, a larger model may display weaker performance. For example, on the DS task, Llama-2-70B exhibits inferior performance compared to Llama-2-13B across all three evaluation metrics. On the CI task, although Yi-34B demonstrates higher accuracy than Yi-6B, it shows higher uncertainty.
@@
C.3 Effects of Instruction Finetuning (Cont.)
Figures 11-13 depict the average results across five tasks of the Yi series, DeepSeek series, and Falcon series, as well as their instruction-finetuned counterparts. Recall that for the instruction-finetuned version, two approaches are adopted to prepare the prompt input. The first approach adheres to the format of the instruction data, which is denoted as Chat-V1. The second approach employs the same prompt format as the base version and is denoted as Chat-V2. For the Yi series, it is observed that both Chat-V1 and Chat-V2 consistently yield inferior performance than the base model concerning all three evaluation metrics. In contrast, for the DeepSeek series, both Chat-V1 and Chat-V2 exhibit enhanced performance in terms of Acc and UAcc. Nevertheless, Chat-V1 results in greater uncertainty compared to the base model for both DeepSeek-7B and DeepSeek-70B. Chat-V2 also demonstrates higher uncertainty relative to the base model for DeepSeek-7B. For the Falcon series, it is observed that both Chat-V1 and Chat-V2 consistently lead to higher uncertainty than the base model, even though Chat-V2 achieves better performance in terms of Acc and UAcc.
@@ C.4 Effects of Error Rate
Conformal prediction ensures that the prediction set encompasses the true label with a probability no less than $1-\alpha$, where $\alpha$ denotes a user-specified error rate. In consideration of this, it is imperative to investigate the impact of varying the value of the error rate $\alpha$ on the prediction set and, subsequently, the estimation of uncertainty. For this purpose, we vary the value of $\alpha$ within the range of 0.05 to 0.3 and report the results of the Llama-2 series in Figure 14. It is observed that as the value of $\alpha$ increases, the coverage rate decreases monotonically. This outcome is anticipated, as a higher error rate implies a reduced probability of the true label being included in the prediction set. Nevertheless, the coverage rate consistently remains greater than $1-\alpha$. This observation reaffirms the statistical guarantee provided by conformal prediction in generating the prediction set. It is also observed that the average set size (SS) decreases monotonically with an increase in the value of $\alpha$. This observation is logical, as a larger error rate suggests that the prediction set can miss the true label with a higher probability and consequently, have a smaller size. In terms of UAcc, according to its definition, it is intuitive that its value will increase as the value of SS decreases. Our results corroborate this expectation. Another noteworthy observation is that, regardless of the value of $\alpha$, Llama-2-70B consistently exhibits lower uncertainty than Llama-2-13B, and Llama-2-13B consistently displays lower uncertainty than Llama-2-7B. This finding is of paramount importance, as it demonstrates that although different values of the error rate $\alpha$ can lead to varying sizes of the prediction set, the relative rankings of different LLMs based on uncertainty remain unchanged.
LLMs Acc (%) $\uparrow$ Coverage Rate (%) SS $\downarrow$ UAcc (%) $\uparrow$
Average Joint Average Joint Average Joint
Yi-34B 81.46 94.22 94.47 1.93 2.18 121.09 112.12
Qwen-72B 78.05 93.33 92.29 2.06 2.10 109.96 102.93
Qwen-14B 74.00 92.94 94.04 2.17 2.39 102.28 86.46
Llama-2-70B 72.24 93.11 92.96 2.16 2.23 94.62 84.42
DeepSeek-67B 71.66 92.21 91.75 2.15 2.29 91.07 79.68
Yi-6B 68.97 92.09 92.77 2.36 2.49 83.53 71.18
Mistral-7B 64.14 91.78 92.05 2.43 2.47 70.84 64.63
Llama-2-13B 60.42 91.60 91.97 2.56 2.65 61.07 56.68
Qwen-7B 59.87 92.07 92.70 2.63 2.69 64.34 56.12
InternLM-7B 49.31 90.96 91.08 3.41 3.45 40.83 35.15
Llama-2-7B 46.53 90.53 90.53 3.09 3.15 38.71 36.22
DeepSeek-7B 45.87 90.48 90.59 3.13 3.18 36.80 35.36
Qwen-1.8B 42.34 90.73 90.60 3.38 3.39 32.95 30.60
Falcon-40B 34.50 90.23 90.15 3.48 3.49 24.85 24.21
MPT-7B 27.18 90.19 90.26 3.57 3.61 18.70 18.47
Falcon-7B 24.84 90.19 90.25 3.75 3.81 16.28 15.95
@@
C.5 Unify All Tasks as One
In our previous analyses, we treat each task as an independent one. Therefore, we need to calculate a separate conformal threshold for each task and generate the prediction set based on the task-specific threshold. However, considering that LLMs are capable of solving multiple tasks, it is possible to consider all five tasks in this study as a single unified task (assuming that the datasets corresponding to the five tasks are drawn from a joint distribution). By doing so, we only need to calculate one conformal threshold and generate the prediction set for all tasks based on this shared threshold. In particular, we combine the calibration sets of all tasks into one and similarly merge the test sets of all tasks into one. The results of various LLMs using this unified approach are presented in Table 5, where we also include the average results of the five tasks when treated individually for comparison purposes. It can be observed that when treating all tasks as a single (joint) one, all LLMs are still able to meet the coverage guarantee requirement. However, they exhibit higher uncertainty in terms of the average set size (SS), which in turn leads to lower performance with respect to UAcc. This finding suggests that while LLMs are indeed capable of addressing multiple tasks, it remains crucial to analyze each task independently since LLMs can demonstrate varying degrees of uncertainty across different tasks.
(a) LAC
(b) APS
(c) LAC + APS (average)
Figure 15: Average results across five tasks of the Llama-2 series when varying the softmax temperature $\tau$.
C.6 Effects of Softmax Temperature
@@
where $z=(z_{1},\dots,z_{m})\in\mathbb{R}^{m}$. A higher temperature results in a more uniform probability distribution (i.e. with higher entropy and "more random"), while a lower temperature leads to a sharper probability distribution, with one value dominating. In this part, we investigate how the temperature $\tau$ affects uncertainty estimation when using conformal prediction. We modify the value of $\tau$ from 0.2 to 2.0 and report the results of the Llama-2 series in Figure 15. Note that the temperature does not affect accuracy, so we omit results regarding accuracy.
@@
It is observed that when using LAC as the conformal score function, we can obtain relatively stable performance in terms of both SS and UAcc. However, APS is highly sensitive to the temperature $\tau$. When $\tau$ takes small values, APS results in high uncertainty, which is somewhat counter-intuitive, as a low temperature leads to sharp distributions and, consequently, low entropy. Intuitively, low entropy suggests low uncertainty. However, when we vary the value of $\tau$, we do not change the LLM’s accuracy. Thus, a low temperature causes the LLM to be overconfident. APS penalizes this phenomenon by producing large prediction sets (implying high uncertainty). This is a desirable property of APS. Nonetheless, we observe that when $\tau$ takes large values, APS provides low uncertainty estimation, which is undesirable. This is because when $\tau$ takes large values, the probability distribution is close to a uniform distribution (indicating high entropy), and the LLM becomes underconfident. Therefore, the uncertainty estimation should be high rather than low. In light of this, we have taken the average value of LAC and APS to achieve a more robust estimation of uncertainty in our analysis.
In summary, these observations suggest that: (I) entropy is not a reliable metric for evaluating uncertainty, (II) LAC is not very sensitive to the softmax temperature $\tau$, (III) a small value of $\tau$ results in the LLM being overconfident, and APS penalizes this by estimating high uncertainty, and (IV) however, APS cannot handle the case where the LLM is underconfident, so it is crucial to take the average results of LAC and APS to achieve a robust quantification of uncertainty.
@@ Appendix D Prompt Templates
{Demonstrations in the same format as the question shown below except that the true answers are provided for questions in the demonstrations} Context/Dialogue/Document: {The context or dialogue history or document corresponding to the following question} Question: {Question} Choices: A. {Content of option A} B. {Content of option B} C. {Content of option C} D. {Content of option D} E. I don’t know F. None of the above Answer:
@@
Context/Dialogue/Document: {The context or dialogue history or document corresponding to the following question} Question: {Question} Choices: A. {Content of option A} B. {Content of option B} C. {Content of option C} D. {Content of option D} E. I don’t know F. None of the above Answer:
@@
(for the QA task) Below are some examples of multiple-choice questions about question answering. Each question should be answered based on your world knowledge and problem solving ability./ (for the RC task) Below are some examples of multiple-choice questions about reading comprehension. Each question should be answered based on the given context and commonsense reasoning when necessary./ (for the CI task) Below are some examples of multiple-choice questions about commonsense natural language inference. For each question, there is a given context and the answer is the option that most likely follows the context./ (for the DRS task) Below are some examples of multiple-choice questions about dialogue response selection. For each question, the answer is the option that represents the most suitable response for the given dialogue history, without hallucination and non-factual information./ (for the DS task) Below are some examples of multiple-choice questions about document summarization. For each question, the answer is the option that accurately summarizes the given document without hallucination and non-factual information.
@@
Context/Dialogue/Document: {The context or dialogue history or document corresponding to the following question} Question: {Question} Choices: A. {Content of option A} B. {Content of option B} C. {Content of option C} D. {Content of option D} E. I don’t know F. None of the above Answer:

### v2

@@ C.1 Detailed Results of LAC and APS
Table 1 and Table 2 have presented the average results derived from the two conformal score functions, namely, LAC and APS. In this part, we analyze the results associated with each conformal score function. The detailed results are reported in Table 6, Table 7, and Table 8.
@@
It is evident that for both conformal score functions, the ranking of LLMs based on accuracy can be different from that based on uncertainty. These results reaffirm the importance of considering uncertainty in order to evaluate the performance of LLMs in a more holistic manner. Another notable observation is the difference in the uncertainty estimations produced by LAC and APS. In general, APS tends to produce larger prediction sets. More importantly, APS can lead to a significantly different ranking of LLMs based on uncertainty compared to LAC. For example, on the CI task, Qwen-14B secures the lowest uncertainty when LAC is utilized as the conformal score function. However, when APS is employed as the conformal score function, Qwen-14B is ranked sixth. This observation suggests that it is essential to average the results of the two conformal score functions to provide a more accurate quantification of uncertainty. Last but not least, both conformal score functions can achieve high coverage rates, verifying again the rationality of relying on prediction set size to estimate uncertainty. It is also noted that APS tends to achieve higher coverage rates than LAC due to its larger prediction sets in most cases.
@@
C.2 Effects of Model Scale (Cont.)
Figures 7-10 illustrate the performance outcomes of the Llama-2 series, Yi series, DeepSeek series, and Falcon series. It is observed that while in general, increasing model size can lead to stronger performance, on some tasks, a larger model may display weaker performance. For example, on the DS task, Llama-2-70B exhibits inferior performance compared to Llama-2-13B across both accuracy and uncertainty. On the CI task, although Yi-34B demonstrates higher accuracy than Yi-6B, it shows higher uncertainty.
C.3 Effects of Instruction Finetuning (Cont.)
Figures 11-13 depict the average results across five tasks of the Yi series, DeepSeek series, and Falcon series, as well as their instruction-finetuned counterparts. Recall that for the instruction-finetuned version, two approaches are adopted to prepare the prompt input. The first approach adheres to the format of the instruction data, which is denoted as Chat-V1. The second approach employs the same prompt format as the base version and is denoted as Chat-V2. For the Yi series, it is observed that both Chat-V1 and Chat-V2 consistently yield inferior performance than the base model in terms of both Acc and SS. In contrast, for the DeepSeek series, both Chat-V1 and Chat-V2 exhibit enhanced performance in terms of Acc. Nevertheless, Chat-V1 results in greater uncertainty compared to the base model for both DeepSeek-7B and DeepSeek-70B. Chat-V2 also demonstrates higher uncertainty relative to the base model for DeepSeek-7B. For the Falcon series, it is observed that both Chat-V1 and Chat-V2 consistently lead to higher uncertainty than the base model, even though Chat-V2 achieves better performance in terms of Acc.
@@ C.4 Effects of Error Rate
Conformal prediction ensures that the prediction set encompasses the true label with a probability no less than $1-\alpha$, where $\alpha$ denotes a user-specified error rate. In consideration of this, it is imperative to investigate the impact of varying the value of the error rate $\alpha$ on the prediction set and, subsequently, the estimation of uncertainty. For this purpose, we vary the value of $\alpha$ within the range of 0.05 to 0.3 and report the results of the Llama-2 series in Figure 14. It is observed that as the value of $\alpha$ increases, the coverage rate decreases monotonically. This outcome is anticipated, as a higher error rate implies a reduced probability of the true label being included in the prediction set. Nevertheless, the coverage rate consistently remains greater than $1-\alpha$. This observation reaffirms the statistical guarantee provided by conformal prediction in generating the prediction set. It is also observed that the average set size (SS) decreases monotonically with an increase in the value of $\alpha$. This observation is logical, as a larger error rate suggests that the prediction set can miss the true label with a higher probability and consequently, have a smaller size. Another noteworthy observation is that, regardless of the value of $\alpha$, Llama-2-70B consistently exhibits lower uncertainty than Llama-2-13B, and Llama-2-13B consistently displays lower uncertainty than Llama-2-7B. This finding is of paramount importance, as it demonstrates that although different values of the error rate $\alpha$ can lead to varying sizes of the prediction set, the relative rankings of different LLMs based on uncertainty remain unchanged.
Figure 15: Average results across five tasks of InternLM-7B when varying the softmax temperature $\tau$. We compare the performance of conformal prediction (CP) and perplexity (PPL).
LLMs Acc (%) $\uparrow$ CR (%) SS $\downarrow$
Average Joint Average Joint
Yi-34B 81.46 94.22 94.47 1.93 2.18
Qwen-72B 78.05 93.33 92.29 2.06 2.10
Qwen-14B 74.00 92.94 94.04 2.17 2.39
Llama-2-70B 72.24 93.11 92.96 2.16 2.23
DeepSeek-67B 71.66 92.21 91.75 2.15 2.29
Yi-6B 68.97 92.09 92.77 2.36 2.49
Gemma-7B 65.74 92.43 92.21 2.38 2.49
Mistral-7B 64.14 91.78 92.05 2.43 2.47
Llama-2-13B 60.42 91.60 91.97 2.56 2.65
Qwen-7B 59.87 92.07 92.70 2.63 2.69
InternLM-7B 49.31 90.96 91.08 3.41 3.45
Llama-2-7B 46.53 90.53 90.53 3.09 3.15
DeepSeek-7B 45.87 90.48 90.59 3.13 3.18
Qwen-1.8B 42.34 90.73 90.60 3.38 3.39
Falcon-40B 34.50 90.23 90.15 3.48 3.49
MPT-7B 27.18 90.19 90.26 3.57 3.61
Falcon-7B 24.84 90.19 90.25 3.75 3.81
@@
C.5 Effects of Softmax Temperature
@@
where $z=(z_{1},\dots,z_{m})\in\mathbb{R}^{m}$. A higher temperature results in a more uniform probability distribution (i.e. with higher entropy and "more random"), while a lower temperature leads to a sharper probability distribution, with one value dominating. In this part, we investigate how the temperature $\tau$ affects uncertainty estimation when using conformal prediction and perplexity, respectively. We modify the value of $\tau$ from 0.2 to 10.0 and report the results of InternLM-7B in Figure 15. Note that the temperature does not affect accuracy, so we omit results regarding accuracy.
It is observed that when using conformal prediction, we can obtain relatively stable performance in terms of both coverage rate and uncertainty (SS). However, perplexity is highly sensitive to the temperature $\tau$. When $\tau$ takes small values, perplexity results in high certainty but low coverage rate. Note that when we vary the value of $\tau$, we do not change the LLM’s accuracy. Thus, a low temperature causes the LLM to be overconfident. In this scenario, it is actually not ideal to estimate low uncertainty. Conformal prediction penalizes this phenomenon by producing large prediction sets (implying high uncertainty), which is more desirable. It is also observed that when $\tau$ takes large values and the probability distribution is close to a uniform distribution (indicating high entropy and that the LLM becomes underconfident), perplexity produces large prediction sets, which are uninformative. In contrast, conformal prediction is able to produce compact prediction sets consistently. This advantage is attributed to the use of a calibration set in conformal prediction. In summary, these observations suggest that conformal prediction is a more reliable method than entropy for evaluating uncertainty.
C.6 Unify All Tasks as One
In our previous analyses, we treat each task as an independent one. Therefore, we need to calculate a separate conformal threshold for each task and generate the prediction set based on the task-specific threshold. However, considering that LLMs are capable of solving multiple tasks, it is possible to consider all five tasks in this study as a single unified task (assuming that the datasets corresponding to the five tasks are drawn from a joint distribution). By doing so, we only need to calculate one conformal threshold and generate the prediction set for all tasks based on this shared threshold. In particular, we combine the calibration sets of all tasks into one and similarly merge the test sets of all tasks into one. The results of various LLMs using this unified approach are presented in Table 9, where we also include the average results of the five tasks when treated individually for comparison purposes. It can be observed that when treating all tasks as a single (joint) one, all LLMs are still able to meet the coverage guarantee requirement. However, they exhibit higher uncertainty in terms of the average set size (SS). This finding suggests that while LLMs are indeed capable of addressing multiple tasks, it remains crucial to analyze each task independently since LLMs can demonstrate varying degrees of uncertainty across different tasks.
@@ Appendix D Prompt Templates
Context/Dialogue/Document: {The context or dialogue history or document corresponding to the following question}
Question: {Question}
Choices:
A.
{Content of option A}
B.
{Content of option B}
C.
{Content of option C}
D.
{Content of option D}
E. I don’t know
F. None of the above
Answer:
@@ F. None of the above
Context/Dialogue/Document: {The context or dialogue history or document corresponding to the following question}
Question: {Question}
Choices:
A. {Content of option A}
B. {Content of option B}
C. {Content of option C}
D. {Content of option D}
E. I don’t know
F. None of the above
Answer:
@@
(for the QA task) Below are some examples of multiple-choice questions about question answering. Each question should be answered based on your world knowledge and problem solving ability./
(for the RC task) Below are some examples of multiple-choice questions about reading comprehension. Each question should be answered based on the given context and commonsense reasoning when necessary./
(for the CI task) Below are some examples of multiple-choice questions about commonsense natural language inference. For each question, there is a given context and the answer is the option that most likely follows the context./
(for the DRS task) Below are some examples of multiple-choice questions about dialogue response selection. For each question, the answer is the option that represents the most suitable response for the given dialogue history, without hallucination and non-factual information./
(for the DS task) Below are some examples of multiple-choice questions about document summarization. For each question, the answer is the option that accurately summarizes the given document without hallucination and non-factual information.
@@
Context/Dialogue/Document: {The context or dialogue history or document corresponding to the following question}
Question: {Question}
Choices:
A. {Content of option A}
B. {Content of option B}
C. {Content of option C}
D. {Content of option D}
E. I don’t know
F. None of the above
Answer: